Determine the profit and overhead markup for a company who wants to maintain a 13.2% gross profit margin.
Answer #1

Solution:

Given:

Gross profit margin=13.2%

Profit and overhead markup of a company is calculating by using following formula:

profit and overhead markup=(Gross profit margin)/(100-(gross profit margin))

=(13.2)/(100-13.2)

=13.2/86.8

=0.15207*100(convert into percentage)

=15.207%

Therefore profit and overhead markup of a company=15.207%